OpenStudy (anonymous):

In Sarah's geometry class, her teacher drew several quadrilaterals on the computer that were projected for the class to see. There were 5 trapezoids, 12 rectangles, 5 squares, and 8 rhombuses. What is the least number of quadrilaterals Sarah's teacher could have drawn? A 30 B 22 C 20 D 17

OpenStudy (anonymous):

You can add 2 squares to create a rectangle, and you can have combinations of rhombi and trapezoids to create squares/rectangles. @praxer

OpenStudy (praxer):

ya, i didn't think of that... :P

OpenStudy (praxer):

In that way, it should be 17, cause, rectangles will add to form another rectangle, since the total is 12, so half will be present, which is 6. Than squares will combine to form rectangles. Since 5 are there. so 4 will combine and reduce to half i.e.2. and 1 square will be alone. than rombi and trapezoids will combine 5, rombi and trapez. will form a sqr/rect. and 3 rombi will be left. so the total sum is, 6+2+1+5+3=6+3+8=9+8=17

OpenStudy (anonymous):

Okay, thanks, I just thought that their was some sort of formula, but i guess you have to just figure it out quickly. This is from a timed test btw

OpenStudy (praxer):